# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

### v2

@@ Abstract
The interpretation of sum-frequency generation (SFG) spectra has been severely limited by the absence of quantitative theoretical predictions of higher-order multipole contributions. The previously unknown magnetic dipole and electric quadrupole contributions are determined by bulk properties but appear in all experimental SFG spectra, obscuring the connection between measured spectra and interfacial structure. We present the simulation-based framework that predicts the full set of multipole spectral contributions. This framework also yields depth-resolved spectra, enabling the precise localization of spectroscopic features. Applied to the air-water interface, our approach achieves quantitative agreement with experimental spectra for different polarization combinations in both the bending and stretching regions. Higher-order multipole contributions are crucial for correctly interpreting SFG spectra: in the bending band, the electric dipole and the magnetic dipole contributions have similar intensities, and the electric quadrupole contribution is significantly larger. In the OH-stretch region, the electric quadrupole contribution is found to be responsible for the mysterious shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$. Crucially, subtracting the quadrupole and magnetic contributions isolates the second-order electric dipole susceptibility, which is a quantitative probe for interfacial orientational anisotropy. This electric-dipole susceptibility reveals a pronounced biaxial ordering of water at the air-water interface. By resolving a fundamental limitation of the interpretation of SFG spectroscopy, our framework allows for the detailed extraction of interfacial water ordering from SFG spectra.
@@ I Introduction
Interfaces are crucial in biological [1, 2, 3, 4, 5] and physicochemical [6, 7, 8, 9, 10] applications. In particular, the air-water interface plays an essential role in catalysis [11, 12, 13, 14, 15], atmospheric [16, 17, 18], and prebiotic [19, 20] chemistry. Undoubtedly, understanding the microscopic structure of interfaces is essential. The inhomogeneous interfacial region is only a few angstroms thick [21], or up to a few nanometers if the interface is charged [22, 23], and thus much thinner than the bulk region. This presents significant challenges in spectroscopy, as signals from the interface must be distinguished from bulk signals. An elegant solution is to measure the second-order susceptibility, which is non-zero only when the spatial inversion symmetry is broken, using sum frequency generation (SFG) spectroscopy. Typically, the bulk medium is isotropic and, therefore, the interface’s structure accounts for a significant part of the SFG spectrum. The interpretation of experimental spectra usually assumes that second-order radiation originates from an interfacial electric dipole layer oscillating at the sum of the applied beams’ frequencies. In this approximation, the SFG signal is determined by a second-order electric dipole susceptibility [24, 25] and model calculations allow for the determination of the orientation of chemical bonds [26, 27, 28], the effective interfacial dielectric constant [29, 30, 31], the position of the dipole layer [29, 30], the surface potential [32, 33, 34, 35], and water structure around macromolecules [36, 37] from experiments. However, this dipole-layer picture neglects higher-order multipole contributions present in the SFG signal. Those multipole contributions to SFG spectra are proportional to higher-order response functions that are non-zero even in isotropic bulk media [38, 39, 40, 41, 42] and cannot be easily experimentally separated from the measured signal. To determine the structure of an interface from SFG spectra, one needs to identify the electric dipole component, which serves as a fingerprint of the interface structure. Thus, accurate theoretical predictions of higher-order multipole contributions are essential. Off-resonant multipole contributions, measured in second-harmonic generation, were predicted in previous works [43, 44]. Resonant multipole SFG contributions where estimated using normal-mode calculations [45] and found to be significant for the water bending band [46]. However, normal-mode methods involve rather drastic approximations, including locality assumptions, and cannot predict the spectral line shape. Experimentally, the multipolar origin of the water bending band has been investigated with SFG at surfactant monolayers [47, 48, 49] and combined SFG/DFG measurements [50], yielding conflicting results. Many studies attribute the bending band to the electric dipole component of the SFG signal [51, 52, 53, 47, 54, 55], in contrast to our findings.
Figure 1: SFG spectra and decomposition into multipole components. (a) & (b): Sketch of the second-order molecular multipole moments that give rise to SFG spectra. Here, $\mu_{i}^{(2)}$, $Q_{ij}^{(2)}$ and $m_{i}^{(2)}$ denote the induced molecular second-order electric dipole, electric quadrupole, and magnetic dipole moments, respectively. The contributions to $\tilde{S}^{(2)}_{yyz}$ are sketched in (a) and to $\tilde{S}^{(2)}_{zzz}$ in (b). The blue region represents water and the white region air. (c): Comparison between predicted and experimental SFG spectra $\tilde{S}^{(2)}_{yyz}$ in the bending frequency region. The SFG spectrum is decomposed into the pure electric dipole ($D$), the electric dipole - electric quadrupole cross (DQ), the electric quadrupole (Q), and the magnetic dipole contribution (M) in (e). (d) & (f): The same analysis for $\tilde{S}^{(2)}_{zzz}$. Results for the OH-stretch frequency region are shown in (g)-(j). Experimental data is taken from Fellows et al. [50], Yu et al. [30] and Chiang et al. [31]. To align with the experimental data, we red-shift our spectra by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in (c)-(f), and by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in (g)-(j). Grey dashed boxes indicate the regions shown in the insets.
@@
Our method, based on time-dependent perturbation theory [56, 57], enables the quantitative prediction of multipolar SFG spectral contributions and constitutes a crucial step toward an accurate determination of the interfacial structure. We compute all relevant multipole contributions, including the magnetic dipole contributions, which were not considered in the past but are found to be sizable and of similar intensity as the electric dipole contribution in the bending band. Following the work of Guyot-Sionnest and Shen [58], we formulate the multipolar second-order electric current density as a response to electric displacement (D) fields perpendicular to the interface and electric (E) fields parallel to the interface, both of which are spatially constant on the relevant length scale. In doing so, we eliminate the need to make ad-hoc assumptions on the interface structure. We observe quantitative agreement between simulated and experimental SFG spectra, with significant multipole contributions. Recent experiments investigated the interfacial dielectric constant [31] and the spatial distribution of anisotropically oriented molecules [29, 59, 21]. Our work reveals that interfacial water exhibits an $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ thick biaxial triple-layer structure, as follows from the second-order electric-dipole susceptibility. This susceptibility is dominated by the biaxial orientational ordering perpendicular to the water molecular dipole axis in the bending region.
@@ II Results
Figure 2: Spatial analysis of the bending band. The mass density profile is shown in (a), the local field factors $f^{\mathrm{VIS}}_{i}(z)$ defined in Equation (1) in (b). Lorentz theory predictions are denoted by dashed horizontal lines: black for vacuum, red and blue for bulk water. For illustrative purposes, a snapshot from the simulation is shown in the background. (c): Comparison of the spatially resolved integrals over the bending band, as defined in Equation (2), between $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and the prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, defined in Equation (33). The electric dipole second-order susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$, defined in Equation (27), is shown at selected positions in (d). These positions are marked in (f), where the corresponding two-dimensional profile is presented. The second-order response profile of the SFG signal $\tilde{s}^{(2)}_{yyz}(z)$ is presented in (e) and $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ in (g). The spectra are red-shifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
@@ II.1 Quantitative Comparison of SFG Spectra with Experiments
In Figure 1, we present SFG spectra $\tilde{S}^{(2)}_{ijk}$ for the polarization combinations yyzyyz and zzzzzz, as defined in Equations (7) and (11). Here, the indices ijkijk denote the polarizations of the second-order electric current density, which produces the observed radiation, the VIS field, and the IR field, respectively. The theoretical framework for predicting multipolar SFG spectra is described in detail in Supplemental Material (SM) Sections I–VI [60]. We observe quantitative agreement between our predictions (red lines) and experiments (black lines) in the bending and stretching bands in Figure 1 (c), (g), (h). We analyze the SFG spectrum by dissecting it into its multipole contributions $\tilde{S}^{(2)}_{ijk}=\tilde{S}^{(2,\mathrm{DD})}_{ijk}+\tilde{S}^{(2,\mathrm{DQ})}_{ijk}+\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$. This decomposition is described in Appendix B. Here, $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ are the pure electric dipole contribution and the electric dipole - electric quadrupole cross contribution, respectively. Importantly $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ describes the interfacial structure, while $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ is created by the linear response of electric dipoles to second-order electric quadrupoles; $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{M})}_{ijk}$ are the electric quadrupole and magnetic dipole contributions, which are in contrast bulk contributions, independent of the interfacial structure. The sum of these contributions $\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$ is called the interfacial quadrupole bulk contribution (IQB) in the literature [41, 42]. To give some intuitive understanding of the physical mechanism producing the different multipolar SFG contributions we schematically illustrate the second-order molecular multipoles of the valence electron charge cloud in Figure 1 (a) and (b). Figure 1 (a) shows the contributions relevant for a y-polarized and (b) for a z-polarized SFG field. The second-order molecular electric dipole moment $\mu_{i}^{(2)}$ is characterized by oscillations of the displacement of the center of the electronic charge distribution and is zero in isotropic media. In contrast, the second-order electric quadrupole moment, characterized by oscillations of the charge distribution width, and the molecular magnetic dipole moment, characterized by oscillations in its angular momentum, as illustrated in Figure 1 (a) and (b), are nonzero in bulk. The second-order electric quadrupole and magnetic dipole moments consist of oppositely directed currents and contribute to the second-order current density via the gradients of their spatial distributions. The bending band SFG spectra in Figure 1 (c)-(f) are dominated by the electric quadrupole contribution (purple), which consists of a positive broad band. The pure electric dipole contribution consists of a negative-positive double peak (Fig. 1 (e) & (f)). The observed lineshape of $\tilde{S}^{(2,\mathrm{DD})}_{yyz}$ qualitatively aligns with previous simulation studies [51, 52, 54], which did not account for higher-order multipole contributions. The magnetic dipole contribution appears only in the yyzyyz spectrum and exhibits an intensity similar to the pure electric dipole contribution (Fig. 1 (e)). The OH-stretch bands are presented in Figure 1 (g)-(j). The pure electric dipole contribution in the stretching region is characterized by a negative–positive double peak, similar to the bending band, as can be seen by comparing the cyan lines in Figure 1 (e) & (i). The difference is that the frequency splitting between the positive and negative components is much larger compared to the spectral linewidth in the OH-stretch region, thus less cancellation takes place and the resulting electric dipole contribution is significantly stronger. The signal arising from the free OH stretch vibrations at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ is largely determined by the pure electric dipole contribution (Fig. 1 (i) & (j)), while significant electric quadrupole contributions to the SFG spectra arise from the more slowly oscillating OH bonds between $3100\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and $3650\text{\,}\mathrm{c}\mathrm{m}^{-1}$. In $\tilde{S}^{(2)}_{yyz}$, the electric quadrupole produces most of the shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ (Fig. 1 (i)). In all spectra presented, $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ is small, and $\tilde{S}^{(2,\mathrm{DQ})}_{yyz}$ is mainly positive, while $\tilde{S}^{(2,\mathrm{DQ})}_{zzz}$ is mainly negative. All spectra have been red-shifted to ensure alignment between theory and experiment. This is reasonable since it has been demonstrated previously that the MB-Pol water model used by us reproduces the experimental vibrational frequencies when nuclear quantum effects are considered [61, 62, 63], while for the linear absorption spectra, it has been demonstrated that nuclear quantum effects do not alter the line shape and absolute intensities much [64, 62].

### v3

@@ Abstract
The interpretation of sum-frequency-generation (SFG) spectra has been severely limited by the absence of quantitative theoretical predictions of higher-order multipole contributions. Magnetic dipole and electric quadrupole contributions are determined by bulk properties but appear in all experimental SFG spectra, obscuring the connection between measured spectra and interfacial structure. We present the simulation-based framework to predict the full set of multipole spectral contributions. This framework also yields depth-resolved spectra, enabling the precise spatial localization of spectroscopic features. Applied to the air-water interface, our approach achieves quantitative agreement with experimental spectra for different polarization combinations in both the bending and stretching regions. Higher-order multipole contributions are crucial for correctly interpreting SFG spectra: in the bending band, the electric dipole and the magnetic dipole contributions have similar intensities, while the electric quadrupole contribution is significantly larger. In the OH-stretch region, the electric quadrupole contribution is found to be in large part responsible for the characteristic shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$. Crucially, subtracting the quadrupole and magnetic contributions isolates the second-order electric dipole susceptibility, which is a quantitative probe for interfacial molecular orientational anisotropy. This electric-dipole susceptibility reveals a pronounced biaxial ordering of water at the air-water interface. By resolving a fundamental limitation of the interpretation of SFG spectroscopy, our framework allows for the detailed extraction of interfacial water ordering from SFG spectra.
@@ I Introduction
Interfaces are crucial in biological [56, 4, 71, 102, 107] and physicochemical [61, 126, 37, 131, 73] applications. In particular, the air-water interface plays an essential role in catalysis [92, 45, 17, 105, 70], atmospheric [143, 8, 74] and prebiotic [33, 51] chemistry. Undoubtedly, understanding the microscopic structure of interfaces is essential. The inhomogeneous interfacial region is only a few angstroms thick [78, 11, 132, 39], or up to a few nanometers if the interface is charged and the salt concentration is small [49, 48], and thus much thinner than the bulk region. This presents significant challenges in spectroscopy, as signals from the interface must be separated from the dominating bulk signals.
@@
An elegant solution is to measure the second-order susceptibility, which is non-zero only when the spatial inversion symmetry is broken, using sum-frequency-generation (SFG) spectroscopy. Typically, the bulk medium is isotropic and, therefore, the interface’s structure accounts for a significant part of the SFG spectrum. The interpretation of experimental spectra usually assumes that second-order radiation originates from an interfacial electric dipole layer oscillating at the sum of the applied beams’ frequencies. In this approximation, the SFG signal is determined by a second-order electric dipole susceptibility [115, 88] and model calculations allow for the determination of the orientation of chemical bonds [144, 133, 123], the effective interfacial dielectric constant [138, 139, 31], the position of the dipole layer [138, 139], the surface potential [46, 96, 100, 119], and water structure around macromolecules [30, 23] from experiments.
However, this dipole-layer picture neglects higher-order multipole contributions present in the SFG signal. Those multipole contributions to SFG spectra are proportional to higher-order response functions that are non-zero even in isotropic bulk media [95, 2, 53, 117, 59] and cannot be easily experimentally separated from the measured signal. To determine the structure of an interface from SFG spectra, one needs to identify the electric dipole component, which serves as a fingerprint of the interface structure. Thus, accurate theoretical predictions of higher-order multipole contributions are essential.
Off-resonant multipole contributions, measured in second-harmonic generation, were predicted in previous works [137, 72]. Resonant multipole SFG contributions were estimated using normal-mode calculations [85] and found to be significant for the water bending band [69]. However, normal-mode methods involve rather drastic approximations, including locality assumptions, and cannot predict the spectral line shape. Experimentally, the multipolar origin of the water bending band has been investigated with SFG at surfactant monolayers [110, 5, 84] and combined SFG and difference-frequency-generation (DFG) measurements [41], yielding conflicting results. Many studies attribute the bending band to the electric dipole component of the SFG signal [90, 93, 34, 110, 112, 111], in contrast to our findings.
[…]
Our method, based on time-dependent perturbation theory [68, 89], enables the quantitative prediction of multipolar SFG spectral contributions and constitutes a crucial step toward an accurate determination of the interfacial structure. We compute all relevant multipole contributions, including the magnetic dipole contributions, which we find to be sizable and of similar intensity as the electric dipole contribution in the bending band.
@@
Following the work of Guyot-Sionnest and Shen [52], we formulate the multipolar second-order electric current density as a response to electric displacement (D) fields perpendicular to the interface and electric (E) fields parallel to the interface, both of which are spatially constant on the relevant length scale. In doing so, we eliminate the need to make ad-hoc assumptions on the interface structure. We observe quantitative agreement between simulated and experimental SFG spectra, with significant multipole contributions.
Recent experiments investigated the interfacial dielectric constant [31] and the spatial distribution of anisotropically oriented molecules [138, 38, 39]. Our work reveals that interfacial water exhibits an $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ thick biaxial triple-layer structure, as follows from the second-order electric-dipole susceptibility. This susceptibility turns out to be dominated by the biaxial orientational ordering perpendicular to the water molecular dipole axis in the bending region.
In Section II.1, we show that quantitative agreement between predicted and experimental SFG spectra is obtained once all relevant multipole contributions are considered. Subtracting all higher-order multipole contribution, the electric-dipole contribution is obtained, which in the bending band is demonstrated to reflect interfacial water orientation in Section II.2. In Section II.3, we show that the interfacial water orientation is significantly biaxial. Sections II.4 and II.5 present the spatially resolved SFG stretch spectrum and the spatially resolved absorption spectrum from the THz to the IR regime, respectively. Finally, we discuss the implications of our findings in Section III and provide a description of our methods in Section IV.
@@ II Results
Figure 2: Depth-dependent analysis of the bending band. The mass density profile is shown in (a), the local field factors in the visual frequency range $f^{\mathrm{VIS}}_{x}(z)$ and $f^{\mathrm{VIS}}_{z}(z)$ defined in Equation (2) in (b). Lorentz theory predictions in bulk water are denoted by dashed horizontal lines. The black solid line denotes the vacuum value. For illustrative purposes, a snapshot from the simulation is shown in the background. (c): Comparison of the spatially resolved integrals over the bending band, as defined in Equation (3), between $\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ and the prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$, defined in Equation (35). The electric dipole second-order susceptibility $\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$, defined in Equation (30), is shown at selected positions in (d). These positions are marked in (f), where the corresponding two-dimensional profile is presented. The second-order response profile of the SFG signal $\tilde{s}^{(2)\prime\prime}_{yyz}(z)$ is presented in (e) and $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$ in (g). The spectra are red-shifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
@@ II.1 Quantitative Comparison of SFG Spectra with Experiments
In Figure 1, we present SFG spectra $\tilde{S}^{(2)}_{ijk}$ for the polarization combinations yyzyyz and zzzzzz, as defined in Equations (8) and (14). Here, the indices ijkijk denote the polarizations of the second-order electric current density, which produces the observed radiation, the VIS field, and the IR field, respectively. The theoretical framework for predicting multipolar SFG spectra is described in detail in Supplemental Information (SI) Sections I–VI.
We observe quantitative agreement between our predictions (red lines) and experiments (black lines) in the bending and stretching bands in Figure 1 (c), (g), (h), (k). We analyze the SFG spectrum by dissecting it into its multipole contributions $\tilde{S}^{(2)}_{ijk}=\tilde{S}^{(2,\mathrm{DD})}_{ijk}+\tilde{S}^{(2,\mathrm{DQ})}_{ijk}+\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$. This decomposition is described in Section IV.3. Here, $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ are the pure electric dipole contribution and the electric dipole - electric quadrupole cross contribution, respectively. Importantly $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ describes the interfacial structure, while $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ is created by the linear response of electric dipoles to fields from second-order electric quadrupoles; $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{M})}_{ijk}$ are the electric quadrupole and magnetic dipole contributions, which are in contrast bulk contributions, independent of the interfacial structure. The sum of these contributions $\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$ is called the interfacial quadrupole bulk contribution (IQB) in the literature [117, 59].
To give some intuitive understanding of the physical mechanisms producing the different multipolar SFG contributions we schematically illustrate the second-order molecular multipoles of the valence electron charge cloud in Figure 1 (a) and (b). Figure 1 (a) shows the contributions relevant for a y-polarized and (b) for a z-polarized SFG field. The second-order molecular electric dipole moment $\mu_{i}^{(2)}$ is characterized by oscillations of the displacement of the center of the electronic charge distribution and is zero in isotropic media. In contrast, the second-order electric quadrupole moment, characterized by oscillations of the charge distribution width, and the molecular magnetic dipole moment, characterized by oscillations in its angular momentum, as illustrated in Figure 1 (a) and (b), are nonzero in bulk. The second-order electric quadrupole and magnetic dipole moments consist of electric currents distributions with vanishing mean and contribute to the second-order current density via the gradients of their spatial distributions.
The bending band imaginary SFG spectra in Figure 1 (c)-(f) are dominated by the electric quadrupole contribution (purple), which consists of a positive broad band, in agreement with previous normal mode calculations [69]. The pure electric dipole contribution consists of a rather weak negative-positive double peak (Fig. 1 (e) & (f)). The observed lineshape of $\tilde{S}^{(2,\mathrm{DD})\prime\prime}_{yyz}$ qualitatively aligns with previous simulation studies [90, 93, 112], which did not account for higher-order multipole contributions. The magnetic dipole contribution appears only in the yyzyyz spectrum and exhibits an intensity similar to the pure electric dipole contribution (Fig. 1 (e)).
The OH-stretch band imaginary SFG spectra are presented in Figure 1 (g)-(j). The pure electric dipole contribution is characterized by a strong negative–positive double peak, in shape similar to the bending band, as can be seen by comparing the cyan lines in Figure 1 (e) & (i). The difference is that the frequency splitting between the positive and negative components is much larger compared to the spectral linewidth in the OH-stretch region, thus less cancellation takes place and the resulting electric dipole contribution is significantly stronger. The positive signal arising from the free OH stretch vibrations at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ is largely determined by the pure electric dipole contribution (Fig. 1 (i) & (j)), while significant electric quadrupole contributions to the SFG spectra arise from the more slowly oscillating OH bonds between $3100\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and $3650\text{\,}\mathrm{c}\mathrm{m}^{-1}$. In $\tilde{S}^{(2)\prime\prime}_{yyz}$, the electric quadrupole produces most of the positive shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ (Fig. 1 (i)). Comparing $\tilde{S}^{(2,\mathrm{DD})\prime\prime}_{yyz}$ (Fig. 1 (i)) and $\tilde{S}^{(2,\mathrm{DD})\prime\prime}_{zzz}$ (Fig. 1 (h)), one sees that $\tilde{S}^{(2,\mathrm{DD})\prime\prime}_{yyz}$ exhibits a small positive shoulder around $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$, whereas $\tilde{S}^{(2,\mathrm{DD})\prime\prime}_{zzz}$ does not. This difference is qualitatively consistent with previous predictions, that were based on electric-dipole approximations [83, 31, 140]. In all spectra presented, $\tilde{S}^{(2,\mathrm{DQ})\prime\prime}_{ijk}$ is small, and $\tilde{S}^{(2,\mathrm{DQ})\prime\prime}_{yyz}$ is mainly positive, while $\tilde{S}^{(2,\mathrm{DQ})\prime\prime}_{zzz}$ is mainly negative.
In Figure 1 (k), we compare several experimental measurements of the absolute SFG spectrum $\big|\tilde{S}^{(2)}_{yyz}\big|=\sqrt{\left(\tilde{S}^{(2)\prime}_{yyz}\right)^{2}+\left(\tilde{S}^{(2)\prime\prime}_{yyz}\right)^{2}}$ [113, 140, 41, 141] with our prediction (red line), showing close agreement within the spread of experimental data across the full frequency range. This comparison also provides an estimate of the experimental uncertainty.
Our predicted spectra have been red-shifted to ensure alignment between theory and experiment. This is reasonable since it has been demonstrated previously that the MB-Pol water model used by us reproduces the experimental vibrational frequencies when nuclear quantum effects are considered [80, 81, 99], while for the linear absorption spectra, it has been demonstrated that nuclear quantum effects do not alter the line shape and absolute intensities much [25, 81]. Quantum correction factors do in fact not appear in spectra to leading order as discussed in the Methods.
Our results reveal significant higher-order multipole contributions. Next, we show how the interfacial electric-dipole contribution in the bending band reflects molecular orientation at the air–water interface.